methanol is burned in a camp stove to provide heat for cooking if 18 g of methanol is burned what will the volume of oxygen measured at 37°C and 98.8 KPA required for complete combustion
1
Expert's answer
2020-12-11T01:46:13-0500

2CH3OH + 3O2 = 2CO2 + 4H2O

n (O2) = 3 x n (CH3OH) /2

M (CH3OH) = 32.04 g/mol

n (CH3OH) = 18 / 32.04 = 0.56 mol

n (O2) = 3 x 0.56 / 2 = 0.84 mol


PV=nRT

V=nRT/P

T = 37+273 = 310 K

R = 8.314 L⋅Pa⋅K−1⋅mol−1

V(O2)=(0.84 x 8.314 x 310) / (98.8 x 1000) = 0.02 L